OpenStudy (anonymous):

Find the real solutions of the equation by graphing. 6x = 9 + x2

OpenStudy (anonymous):

Is it \[6x=9+x ^{^{2}}\]

OpenStudy (anonymous):

thats the question....

OpenStudy (anonymous):

6x=9+x^(2) -x^(2)+6x=9 -x^(2)+6x-9=0 x^(2)-6x+9=0 (x-3)(x-3)=0 x-3=0 x=3

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Poooop

OpenStudy (anonymous):

looks something like that. so the answer is 3